OpenStudy (anonymous):

Multiply and simplify. Assume that all variables represent nonnegative numbers. (-4square root 5)(-2square root 3)

OpenStudy (anonymous):

\[(-4\sqrt{5})(-2\sqrt{3)}\]

OpenStudy (mertsj):

Multiply the coefficients and multiply the radicands.

OpenStudy (anonymous):

square root 15*8

OpenStudy (mertsj):

\[8\sqrt{15}\]

OpenStudy (anonymous):

Then what do I need to do

OpenStudy (mertsj):

you are done.

OpenStudy (anonymous):

thanks